OpenStudy (anonymous):

Find the equation for the normal line to f at the point (-2, f(-2)). f(x) = 3x- x^2

OpenStudy (precal):

normal line means that it is perpendicular

OpenStudy (anonymous):

so what is the (-2, f(-2) for

OpenStudy (anonymous):

i know normal means perpendicular. idk how to get he equation

OpenStudy (precal):

sub x=-2 into the f(x) equation to give you the y value they want you to use. Take the derivative of f(x) to get the slope. Use that slope and that point to find the normal line.

OpenStudy (amistre64):

perp slope are flip and negate

OpenStudy (anonymous):

for teh der, i got f' is 3x - 2x....wat is teh der..i think its wrong

OpenStudy (amistre64):

f'=[3x- x^2]' =3- 2x

OpenStudy (amistre64):

this tells us the slope of the tangent at any point

OpenStudy (amistre64):

-1/f' = slope of the normal at any given point then

OpenStudy (anonymous):

so do i pug in -2 to get teh der now ...is it 7

OpenStudy (amistre64):

since it wants the normal at -2, then yes; use -2

OpenStudy (amistre64):

so we use this slope of "1" in our equation for the line; do you recall the point slope equation for a line?

OpenStudy (anonymous):

wait so the slope is 7 but its supposed to be neg. reciprocl so shudnt it be 1/7 te hslope

OpenStudy (amistre64):

the slope aint 7, its -1

OpenStudy (amistre64):

your right... 7 is good, I was solving some make believe problem in my head lol

OpenStudy (amistre64):

-1/7 is the normal slope yes, well stick with that :)

OpenStudy (anonymous):

my final equation is y +10 = -1/7 x +2/7...is taht correct

OpenStudy (precal):

I stand corrected; for the normal slope use the opposite reciprocal.

OpenStudy (amistre64):

maybe, lets see if i can stick with it :) y = (-1/7)x +(1(-2)/7) + f(-2) y = (-1/7)x -2/7 -10(7)/7 y = (-1/7)x +(-2-70)/7 y = (-1/7)x -72/7 might be good

OpenStudy (amistre64):

y = m(x-Px)+Py y = mx-mPx+Py

OpenStudy (anonymous):

thank u very much :)